The gem emerald is a pure crystal of the mineral beryl, Be3Al2(SiO3)6 How many atoms of beryllium would be in 134.35 grams of beryl?


1
Expert's answer
2021-10-10T10:20:06-0400

M(Be3Al2(SiO3)6) = 537.50 g/mol

n(Be3Al2(SiO3)6) "= \frac{134.35}{537.50}= 0.2499 \;mol"

Looking at the formula of beryl we can see that there are 3 moles of Beryllium in beryl so 1 one mole of beryl gives 3 moles of beryllium, 0.2499 moles of beryl give "0.2499 \times 3 = 0.7498" moles of beryllium. One mole contains 6.022×1023 elementary entities of the substance.

The number of atoms of beryllium "= 0.7498 \times 6.022 \times 10^{23} = 4.515 \times 10^{23} \;atoms"